OpenStudy (anonymous):

If one measurement of a golden rectangle is 3.2 inches, which could be the other measurement? If one measurement of a golden rectangle is 3.2 inches, which could be the other measurement? a. 5.2 inches c. 4.4 inches b. 1.9 inches d. 3.0 inches

Directrix (directrix):

The ratio of length to width of a golden rectangle is the golden ratio (approx) 1.618. Solve 3.2/x = 1/1.618 where x is the other possible measurement.

Directrix (directrix):

I think the argument could also be given for a second possible measure: 3.2/x = 1.618/1 x = ?